RC meeting and Run Plan, Thursday March 9th
Return to the RC Meeting Calendar
Contents
RC Report
- RC: Holly
Summary
- RC Daily report: https://logbooks.jlab.org/entry/4145924
Beam was restored during swing shift yesterday thanks to Jiwan and Zeke. We did note that when the carbon hole checkout gave confusing results, initializing the target ladder gave us a clear carbon hole image (as consistent with Bill's recommendation during the RC meeting yesterday). We lost beam due to accelerator issues for about 2 hours last night, and then restored beam before midnight. Owl shift continued with the slow current ramp-up on the pol-3He cell and hit 45uA this morning. We are back to 45uA production. Bill and Hunter evaluated the target performance this afternoon (https://logbooks.jlab.org/entry/4146074). They concluded that the position of the target as it polarizes seems important (better polarization in No Target position vs Carbon Hole, for example).
Swing summary: https://logbooks.jlab.org/entry/4145384 Owl summary: https://logbooks.jlab.org/entry/4145647 Day summary: https://logbooks.jlab.org/entry/4145914
Recent Issues
- Note that initializing target ladder gave more consistent beam hole analysis: https://logbooks.jlab.org/entry/4145507 (before https://logbooks.jlab.org/entry/4145475)
- Parity DAQ was off for 2 hours this morning: https://logbooks.jlab.org/entry/4145911
- SBS GEMs disabled from replay right now https://logbooks.jlab.org/entry/4146072
- getCharge script doesn't work on runs that ended because the DAQ crashed...
Current Status
- Latest carbon hole image: https://logbooks.jlab.org/entry/4145626
- Running 45uA on pol-3He
Run Plan
- Production on 3He!
- GENII-3Stream config
- PS1=2, PS3=0, PS5=6
- Useful guide from Bill on rates to compare when bringing beam back: https://logbooks.jlab.org/entry/4144886
Data Collection and Polarization Status
- Currently 41%: https://logbooks.jlab.org/entry/4146044
- Data status at 4th-pass: 45.4 C (uncorrected)